# Jarne Verhaeghe IDLab

### v1

@@ 5.2 Implementation
In the case of synthetic data, the true propensity distribution, also known as the oracle distribution, is available. However, in real-world applications, the true propensity distribution is mostly unknown. As a result, any method that relies on propensity is evaluated using both the oracle propensity distribution and an estimated propensity distribution in the experiments, denoted as "Oracle" and "Propensity" in the results respectively. The latter can be approximated by leveraging conformal prediction, specifically CPS. Do note that CPS quantifies total uncertainty and thus also includes the epistemic uncertainty while ideally only the aleatoric uncertainty is included. Additionally, this propensity distribution estimate is not completely guaranteed to be equal to the true conditional propensity distribution, which we theoretically need to get complete finite sample guarantees of validity. Although, in practice, this can still be a valid approximation. A learner is trained on the covariates $X$ to predict the treatment assignment $T$, deemed the propensity learner. Subsequently, a CPS is calibrated for this learner using the calibration set as it is more practical to extract an empirical density distribution compared to standard conformal prediction. This CPS produces an empirical density distribution being a sum of Dirac delta distribution similar to $F_{R}$, thus we require the use of kernel density estimation (KDE) to extract a continuous propensity density function for a treatment value $t$, given covariates $X_{i}$. Do note that KDE interpolates the density and depending on the KDE parameters may introduce additional epistemic error, which is a drawback of estimating the propensity in this manner. The implementation for the propensity estimation is shown in Appendix B.
For the evaluation, several methods were tested and compared, including Gaussian Process, CatBoost with Uncertainty (Duan et al., 2019), Standard Conformal Prediction, Locally Weighted Conformal Prediction (WCP Local), Global Propensity-Weighted Conformal Prediction (WCP Global Oracle and WCP Global Propensity), and Local Propensity-Weighted Conformal Prediction (WCP Local Oracle and WCP Local Propensity). The Gaussian Process was included in the comparison due to its widespread use for UQ in regression problems assuming a normal error distribution (Fiedler et al., 2021). All other approaches were based on the CatBoost model, chosen for its strong out-of-the-box performance (Dorogush et al., 2018). As a result, the "CatBoost with Uncertainty" method was incorporated as a baseline for comparison of UQ.
@@
The propensity learner employed in the propensity-weighted approaches was a CatboostRegressor with 4000 iterations and default hyperparameters. Similarly, the CADRF models were based on CatBoost with 5000 iterations and default hyperparameters. The CatBoost with Uncertainty approach used the same underlying CatBoost model as the CADRF methods to ensure consistency. For the locally weighted conformal approaches, a Gaussian kernel (Theodoridis, 2015) was employed to represent local coverage. The bandwidth parameter for the kernel was set as $h=2\cdot(0.2\cdot\sigma_{\hat{\pi}})^{2}$, where $\sigma_{\hat{\pi}}$ denotes the standard deviation of the estimated propensity distribution [^2] The code of the proposed methodology and the experiments are available open-source at https://github.com/predict-idlab/dose-response-conformal-prediction.
5.3 Results
Figure 1 presents the coverage bar plots across all methods for Setup 3 Scenario 1 on the test set. Evaluations on all other setups and scenarios can be found in Appendix D. The bar plots in Figure 1 clearly illustrate the impact of covariate shift in the treatment on coverage guarantees for methods that did not account for this shift. All propensity-weighting methods assumed uniform treatment sampling during evaluation, mimicking the interpretation of a dose-response curve for decision-making for all treatment values, keeping their coverage guarantees.
As can be seen in Figure 1, the global propensity-weighting method shows a high variance in coverage across different experiments. This variance arises due to the calibration process, which considers all possible treatment values between $t_{L}$ and $t_{U}$, including those with minimal or no overlap. Depending on the calibration and test set split, certain samples may receive a significantly large likelihood ratio, thereby assigning considerable weight to those values according to Equation 12. This inflates the size of the prediction intervals, leading to conservative estimates. The oracle estimates are also notably more conservative, as they tend to provide narrower propensity distributions. This increases the frequency of large likelihood ratios when compared to the estimated propensity distribution, where the epistemic uncertainty of the propensity learner is also taken into account by the CPS procedure. On the contrary, for a new sample, the local propensity method uses calibration samples with treatment values close to the predefined value $t_{0}$ and weighting the propensities as well. Our presented approach uses more comparable calibration samples rather than the entire dataset, resulting in more conditional prediction intervals, provided there are enough calibration samples. Our method thus combines the strengths of both the local and the propensity weighting techniques. These trends are further supported in Figure 2, which shows the prediction intervals for all weighting methods alongside the treatment assignment distribution for a specific test observation. This example highlights the necessity of the uniform treatment sampling assumption for the evaluation of dose-response curves, as both the local weighting method and standard conformal prediction produce inaccurate prediction intervals in regions with low treatment overlap. In these regions, there is insufficient data to support predictions for the model, making these predictions unreliable. Consequently, propensity-weighted methods produce much larger prediction intervals in these areas to compensate for this lack of data support. If there is almost no support or extremely low propensity values, then the propensity-weighted methods provide intervals with an infinite width to show that there is no support in these regions. It is important to note, however, that these intervals may be overly conservative if the model has indeed generalized effectively in such regions. The only way to validate this is through additional data collection in these areas to confirm the model’s performance.
@@ 5.3 Results
Figure 1: Barplot of the mean coverage calculated over 40 treatment values in 50 experiments for setup 3 scenario 1. Black dotted line is the ideal coverage.
@@
Note that Schröder et al. (2024) also introduced a conformal prediction method to provide prediction intervals in the continuous treatment setting. However, we did not include a direct comparison in this study due to the high computational complexity of their approach, which would require several years to complete the same experiments we executed in a matter of hours. For a more detailed comparison, including a discussion of the difference in assumptions and methodologies, see Appendix C.
@@
Implementing local propensity weighting in practice is less straightforward as it involves calibrating for a set of predefined treatment values and either storing these models for later use during inference or performing this action in parallel. This has the advantage that it allows conditional prediction intervals to be calculated more quickly during inference. However, a drawback is that evaluating a treatment value not included in the predefined set requires recalibration, and must be considered for inference. Still, this approach is particularly useful in fields like drug dosing, where treatment ranges are often predefined and personalized CADRF is highly relevant or where inference of new treatment values is not time-critical. Additionally, an important factor to consider is the effective sample size $\hat{n}$ in local propensity weighting (Tibshirani et al., 2019; Jonkers et al., 2024a). Reweighting $F_{R}(y)$ can significantly reduce the effective sample size, which increases variability in empirical coverage compared to standard conformal prediction. This issue is especially pronounced in regions with low treatment overlap, where the effective sample size can become extremely small. However, as prediction intervals with infinite length are possible using weighted conformal prediction, these infinite intervals additionally provide information to the user where the model cannot be trusted adding an interpretability layer to the UQ. In the current work, only an S-learner was used as a CADRF estimator which could influence the epistemic error, so in future work, more specialised dose-response models can be used to reduce the interval widths and provide even more informative prediction intervals.
@@
Our current approach can be readily extended by incorporating other conformal prediction frameworks that support weighted conformal prediction, such as adaptive conformal prediction (Romano et al., 2019) or weighted conformal predictive systems (Jonkers et al., 2024a). Additionally, the weighting can be further expanded or changed to account for other types of covariate shifts in a similar manner or serve different purposes such as evaluating interventions of causal effects, thus broadening the applicability of the proposed method.
@@ 6 Conclusion
Acknowledgments

### v2

@@ 5.3.2 Implementation
In the case of synthetic data, the true propensity distribution, also known as the oracle distribution, is available. However, in real-world applications, the true propensity distribution is mostly unknown. As a result, any method that relies on propensity is evaluated using both the oracle propensity distribution and an estimated propensity distribution in the experiments, denoted as ”Oracle” and ”Propensity” in the results respectively. The estimated distribution in this work is obtained using the Conformal Prediction System (CPS), leveraging conformal prediction, though other propensity estimators could also be used. Do note that CPS quantifies total uncertainty and thus also includes the epistemic uncertainty while ideally only the aleatoric uncertainty is included. Additionally, this propensity distribution estimate is not completely guaranteed to be equal to the true conditional propensity distribution, which we theoretically need to get complete finite sample guarantees of validity. Although, in practice, this can still be a valid approximation. A learner is trained on the covariates $X$ to predict the treatment assignment $T$, deemed the propensity learner. Subsequently, a CPS is calibrated for this learner using the calibration set as it is more practical to extract an empirical density distribution compared to standard conformal prediction. Since CPS produces an empirical density distribution being a sum of Dirac delta distribution similar to $F_{R}$, kernel density estimation (KDE) is applied to derive a continuous propensity density function for a treatment value $t$, given covariates $X_{i}$. Do note that KDE interpolates the density and depending on the KDE parameters may introduce additional epistemic error, which is a drawback of estimating the propensity in this manner. The implementation and computational discussion for Global and Local Propensity WCP is presented in Appendix B.1 and our propensity estimation in Appendix B.2.
@@
For the evaluation, several baseline methods were tested and compared, including Gaussian Process, CatBoost with Uncertainty (duan_ngboost_2019), Standard Conformal Prediction, and Locally Weighted Conformal Prediction (WCP Local, using weights $w_{l}$). The assumed target interventional distribution is a uniform distribution, to mimic that every dose could be equally likely queried. For the proposed propensity methods, we included both variations, using their respective weights: Global Propensity-Weighted Conformal Prediction (WCP Global Oracle and WCP Global Propensity using $w_{g,p}$) and Local Propensity-Weighted Conformal Prediction (WCP Local Oracle and WCP Local Propensity, using $w_{l,p}$). The Gaussian Process was included in the comparison due to its widespread use for UQ in regression problems assuming a normal error distribution (fiedler_practical_2021). All other approaches used a CatBoost model for the base CADRF learner, chosen for its strong out-of-the-box performance (dorogush_catboost_2018). As a result, the ”CatBoost with Uncertainty” method was incorporated as a baseline for comparison of UQ. The propensity learner employed in the propensity-weighted approaches was a CatboostRegressor with 4000 iterations and default hyperparameters. Similarly, the CADRF models were a CatBoost model with 5000 iterations and default hyperparameters. The CatBoost with Uncertainty approach used the same underlying CatBoost model as the other methods to ensure consistency. For the locally weighted conformal approaches, a Gaussian kernel (theodoridis_chapter_2015) was employed to represent local coverage. The bandwidth parameter for the kernel was set as $h=2\cdot(0.2\cdot\sigma_{\hat{\pi}})^{2}$, where $\sigma_{\hat{\pi}}$ denotes the standard deviation of the estimated propensity distribution. This bandwidth parameter was originally selected using 5-fold cross-validation for every scenario and evaluating the calibration on a held-out set. However, we found that $h=2\cdot(0.2\cdot\sigma_{\hat{\pi}})^{2}$ served as a robust heuristic that provided sufficient and consistent performance across all experiments and scenarios. We note that the actual performance and tightness of the prediction intervals can be further improved by selecting an optimal, use-case-specific bandwidth parameter through cross-validation  [^2] The code of the proposed methodology and the experiments are available open-source at https://github.com/predict-idlab/dose-response-conformal-prediction.
5.4 Semi-synthetic
In addition to the fully synthetic dataset, we evaluate the proposed method on a semi-synthetic dataset derived from AMICAS, an open-source patient simulator for anaesthesia drug administration, designed for multi-drug dosing control in surgical patients 9316655. We used this simulator to simulate the bispectral index (BIS) of the patients, a commonly used measure of anaesthetic depth. The advantage of using this simulator is that we can also query the counterfactual using the simulator and thus evaluate the methods on the ground truth. This dataset consists of 1000 randomly generated patients with physiologically plausible characteristics sampled from the following distributions with a slightly biased dataset having more males (Gender = 0):
[…]
$BMI$ $=\frac{Weight}{(Height/100)^{2}}$
[…]
Each patient was administered a dose of Propofol $T_{propofol}$, an intravenous anaesthetic, drawn from a normal distribution representing the conditional propensity distribution incorporating confounding effects from body mass index (BMI) and lean body mass (LBM):
$Propofol\sim\text{Normal}\left(\mu_{prop}=0.05+\frac{0.0005\ age^{2}+0.3\frac{BMI}{40}^{3}+0.4\ e^{\frac{LBM}{40}}}{25},\sigma=0.1\right)$ (24)
$T_{Propofol}=max(0.05,Propofol)$
Due to simulator constraints, a minimum propofol dose of 0.05 was enforced. Additionally, we incorporated three commonly co-administered anaesthesia-related drugs: Atracurium, Dopamine, and Sodium Nitroprusside (SNP). These can influence the effect of propofol on the BIS and are sampled as follows:
[…]
Given these patient characteristics and drug administrations, the AMICAS simulator then calculated the BIS ($BIS_{AMICAS}$) using a Minto pharmacokinetic model, an anesthesiologist-in-the-loop setting, and no disturbance profile. The AMICAS simulator simulates the BIS across time, starting at 0 minutes to a maximum of $300$ minutes representing the simulation times.
Since the $BIS_{AMICAS}$ values are derived from a simulation rather than real-world measurements, we introduced heteroscedastic noise to better approximate more complex uncertainty which can occur in reality. This noise accounts for both age-dependent measurement variability and increased uncertainty when the administered propofol deviates from the expected dose for a given patient. Specifically, the noise follows a normal distribution with a standard deviation modulated by age and dose deviation from the mean conditional propensity $\mu_{prop}$ for that patient (see Equation 24):
$\epsilon\sim\text{Normal}\left(0,\left(\frac{|T-\mu_{prop}|}{0.1}\right)^{2}*\left((1-I(age<69))*3+(I(age<69))*6\right)\right)$ (25)
$BIS=BIS_{AMICAS}+\epsilon$ (26)
This standard deviation increases variability in the BIS values for older patients and for cases where the administered dose substantially deviates from the expected treatment value, introducing additional confounding in the form of added noise. This noise is also indedepently sampled for every time point and every patient.
The experimental evaluation largely follows the synthetic benchmarking with some modifications: Model training was performed using 1000 iterations for all CatBoost models and every model is trained to predict the BIS given the following features: time, administered propofol $T_{propofol}$, co-administered medication (Atracurium, Dopamine, and SNP), age, gender, LBM, BMI, height, and weight. The dataset was split on patient IDs into training, calibration, and test sets with proportions of $53.3\%$, $26.6\%$, and $20\%$, respectively. The significance levels were also evaluated at $\alpha=[0.1,0.05]$ and the bandwidth of the KDE is set to $0.02$ instead of $1$. We included $12$ different time points in each model, with $10$ minutes being the first BIS measurement: $10,30,60,80,100,120,150,170,200,220,250,$ and $300$ minutes resulting in a complete semi-synthetic dataset of $12000$ samples. All models were included in the evaluation except for Gaussian Processes (GP) due to the computational constraints of Gaussian Processes. The propofol treatments $t_{0}$ of test patients were assessed over a dose range from $0.05$ to $0.5$, with increments of $0.01$ resulting in $45$ evaluations per time point per patient. The counterfactual is evaluated using the AMICAS simulator and a new noise value is generated to ensure independent noise across counterfactuals. This is repeated using 100 different random seeds for splitting aggregated, similar to the synthetic benchmarking [^3] The code and data are also available on GitHub.
6 Results and discussion
@@
Figure 1: Barplot of the mean coverage calculated over 40 treatment values in 50 experiments for setup 3 scenario 1. The black dotted line is the ideal coverage.
(a) $\alpha=0.1$
(b) $\alpha=0.05$
Figure 2: Barplot of the mean coverage calculated over 45 treatment values in 100 experiments for the AMICAS semi-synthetic evaluation. The black dotted line is the ideal coverage.
Figure 1 presents the coverage bar plots across all methods for Setup 3 Scenario 1 on the test set. More evaluations, box plots showing calibration on dose-level, and CADRF RMSE on all synthetic setups and scenarios can be found in Appendix D. The bar plots in Figure 1 clearly illustrate the impact of covariate shift in the treatment on coverage guarantees for methods that did not account for this shift.
As can be seen in Figure 1, the global propensity-weighting method shows conservative estimates across different experiments. These conservative estimates arise due to the calibration process, which considers all possible treatment values equally between $t_{L}$ and $t_{U}$, including those with minimal or no overlap. Depending on the calibration and test set split, certain samples may receive a significantly large likelihood ratio, thereby assigning considerable weight to those values according to Equation 12. This inflates the size of the prediction intervals, leading to conservative estimates. The oracle estimates are also notably more conservative, as they tend to provide narrower propensity distributions. This increases the frequency of large likelihood ratios when compared to the estimated propensity distribution, where the epistemic uncertainty of the propensity learner is also taken into account by the CPS procedure. On the contrary, for a new sample, the local propensity method uses calibration samples with treatment values close to the predefined value $t_{0}$ and weighting the propensities as well. Our presented approach uses more comparable calibration samples rather than the entire dataset, resulting in more conditional prediction intervals, provided there are enough calibration samples. Our method thus combines the strengths of both the local and the propensity weighting techniques. The same conclusions for the local propensity calibration can be made on the results of the semi-synthetic benchmarking, presented in Figure 2(b). In the semi-synthetic benchmarking, the global (oracle) model has a high variance in performance attributed to the severe heteroscedastic noise in the semi-synthetic dataset that becomes worse in regions with less treatment overlap.
These trends are additionally supported by Figure 3, which visually shows the prediction intervals for all weighting methods alongside the treatment assignment distribution for a specific test observation in the synthetic dataset Setup 3 Scenario 1. This example highlights the necessity of the uniform treatment sampling assumption for the evaluation of dose-response curves, as both the local weighting method and standard conformal prediction produce inaccurate prediction intervals in regions with low treatment overlap. In these regions, there is insufficient data to support predictions for the model, making these predictions unreliable. Consequently, propensity-weighted methods produce much larger prediction intervals in these areas to compensate for this lack of data support. If there is almost no support or extremely low propensity values, then the propensity-weighted methods provide intervals with an infinite width to show that there is no support in these regions. It is important to note, however, that these intervals may be overly conservative if the model has indeed generalized effectively in such regions. The only way to validate this is through additional data collection in these areas to confirm the model’s performance.
Note that schroder_conformal_2024 also introduced a conformal prediction method to provide prediction intervals in the continuous treatment setting. However, we did not include a direct comparison in this study due to the high computational complexity of their approach, which would require several years to complete the same experiments we executed in a matter of hours. For a more detailed comparison, including a discussion of the difference in assumptions and methodologies, see Appendix C.
@@
Implementing local propensity weighting in practice is less straightforward as it involves calibrating for a set of predefined treatment values and either storing these models for later use during inference or performing this action in parallel. This has the advantage that it allows conditional prediction intervals to be calculated more quickly during inference. However, a drawback is that evaluating a treatment value not included in the predefined set requires recalibration, and must be considered for inference. Consequently, coverage guarantee is not simultaneous over all possible $t_{i}$ if calibrated for a single dose $t$, unless multiple predefined treatment values have been calibrated. Still, this approach is particularly useful in fields like drug dosing, where treatment ranges are often predefined and personalized CADRF is highly relevant, or where inference of new treatment values is not time-critical. Additionally, an important factor to consider is the effective sample size $\hat{n}$ in local propensity weighting (tibshirani_conformal_2019; jonkers_conformal_pred_shift_2024). Reweighting $F_{R}(y)$ can significantly reduce the effective sample size, which increases variability in empirical coverage compared to standard conformal prediction. This issue is especially pronounced in regions with low treatment overlap, where the effective sample size can become extremely small. However, as prediction intervals with infinite length are possible using weighted conformal prediction, these infinite intervals additionally provide information to the user where the model cannot be trusted, adding an interpretability layer to the UQ. In the current work, only an S-learner was used as a CADRF estimator, which could influence the epistemic error, so in future work, more specialised dose-response models can be used to reduce the interval widths and provide even more informative prediction intervals.
7 Extensions and applications
@@
Our current approach can be readily extended by incorporating other conformal prediction frameworks that support weighted conformal prediction, such as adaptive conformal prediction (romano_conformalized_2019) or weighted conformal predictive systems (jonkers_conformal_pred_shift_2024). Additionally, the weighting can be further expanded or changed to account for other types of covariate shifts in a similar manner or serve different purposes, such as evaluating interventions of causal effects, thus broadening the applicability of the proposed method.
The paper’s current setup assumes no covariate shift in the features $X$ between the training, calibration, and test set, i.e., $P_{X}=\tilde{P}_{X}$, to simplify the derivation of the propensity-based weights. However, in real-world applications, covariate shifts are much more common and can hamper the coverage guarantee of conformal prediction, and also thus our proposed method (tibshirani_conformal_2019). If we assume $P_{X}\neq\tilde{P}_{X}$ in equation 15, we observe that this results in adding a multiplicative term that represents the likelihood term for the covariate shift in $X$. As such, both $w_{l,p}^{t}$ and $w_{g,p}$ can be easily adjusted to cover a covariate shift in the test set if the covariate shift is known or can be calculated, analogous to tibshirani_conformal_2019, resulting in the following new weights:
$w_{g,p}(X_{i},T_{i})\propto\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})}{\pi(T_{i}|X_{i})}\frac{d\tilde{P}_{X}}{d\tilde{P}_{X}}$ (27)
and
$w_{l,p}^{t}(X_{i},T_{i})\propto\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})K\left(\frac{T_{i}-t}{h}\right)}{\pi(T_{i}|X_{i})}\frac{d\tilde{P}_{X}}{d\tilde{P}_{X}}$ (28)
Furthermore, because the method is built using conformal prediction, the whole approach is model-agnostic. As such, any possible CADRF model that provides a dose-response curve given features and treatment can be used and thus is not limited to the presented CADRF approach in this paper.
The classic application is in drug dosing, where the goal is to construct a dose-response curve for every individual to facilitate decision-making when determining an optimal dose for a new patient. In a clinical trial, especially phase 1 and phase 2, where the optimal dose is being determined, the weighted conformal dose-response curve can also act as a tool to analyse the results individually while having an estimate of the uncertainty estimates that is not biased by the treatment assignment distribution. It quantifies uncertainty for individual predictions, compensating for any treatment distribution bias. Furthermore, it highlights areas with insufficient data support with infinite prediction intervals, guiding decisions about whether further trials or treatments are necessary for specific patient subgroups. In the regions where there is support, the model predictions provide the CADRF estimate for this patient, and the uncertainty regions show how the outcome would vary.
Treatment is not limited to healthcare. Treatment can be generalized as any intervention or action that opens applications in other domains. For example, in predictive maintenance, the model can optimize decisions by estimating the effect of operating pressure on the remaining useful life of equipment like valves. Similarly, in sales, it can help determine the ideal discount for specific clients to maximize the sold units, demonstrating flexibility in various domains.
Figure 4: A Ceteris Paribus curve generated with Local Propensity WCP.
The application potential is also not limited to actual treatments and interventions. The method can also be used for the explainability of a model. Suppose we fitted a regression model, regressing $X=[X_{1},...,X_{m}]$ on $Y$. $X$ is observed data; thus, any feature can be confounded or biased. By considering a feature $X_{i}$ as a treatment, we can apply the proposed method on this feature instead of a treatment variable, which then provides uncertainty quantification on a Ceteris Paribus curve of a model in a similar manner to a dose-response curve  [^4] A Ceteris Paribus curve visualizes a model’s predictions while keeping all features constant except for one explanatory variable. The x-axis represents the explanatory variable, and the y-axis shows the corresponding predictions.. This curve can then give uncertainty estimates of the ”true” outcome for an individual sample if that sample would have had other values for this particular feature.
An example is shown in Figure 4 using Local Propensity WCP. This example is generated using the Boston Housing data available native in sklearn (sklearn_api), split into a training and calibration set using a $75/25$ split. A CatBoostRegressor using 300 iterations is fitted on a training set, and a propensity CatBoostRegressor with the same number of iterations is fitted on the training set. A CPS is used and calibrated on the calibration set for the propensity distribution estimate, similar to the experimental setup in this paper. No hyperparameter tuning is applied for simplicity, so note that the epistemic uncertainty could be further reduced. The chosen feature for generating a ceteris paribus curve is MedInc, the median income, an important variable in predicting the median house value in this dataset. The figure is for a single data sample where all other variables of this sample are kept constant except for our ”treatment” MedInc. In Figure 4, it is apparent that the prediction intervals go to infinity for MedInc values below $1$ and above $6.5$. This indicates that there is insufficient overlap to evaluate this sample for these values of MedInc, clearly showing a bias in the data distribution of MedInc, given the other features. Consequently, the predictions for a sample with these features but with a MedInc of, e.g., $8$ cannot be trusted as the model is simply doing an interpolation in an out-of-bounds region. In the regions with support, i.e., around $1.5<MedInc<6.5$, we see that the model shows a linear relation with the median house value with relatively small uncertainty bounds. This analysis can be done for any other regression model in a likewise manner.
@@ 8 Conclusion
Statements and Declarations
Funding and conflict of interests